OpenStudy (anonymous):

A novice golfer on the green takes three strokes to sink the ball. The successive displacements are 3.80 m to the north, 1.80 m northeast, and 0.90 m at 30.0° west of south. Starting at the same initial point, an expert golfer could make the hole in what single displacement?

OpenStudy (anonymous):

@ryanvarghese12790

OpenStudy (anonymous):

Horizontal___________Vertical 0___________________3.8 1.8 cos 45°__________1.8 sin45° - 0.9 sin 30°_________-0.9 cos30° 0.823________________4.29 d = 0.823 i + 4.29 j d = 4.37 tan Ө = 4.29 / 0.823 Ө = 79.1° Single displacement is 4.37 m at angle 79.1 ° N of E

OpenStudy (anonymous):

the choices does not fit the answer

OpenStudy (anonymous):

@ryanvarghese12790

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

@ParthKohli

Parth (parthkohli):

You know what vectors are though, right?

OpenStudy (anonymous):



OpenStudy (anonymous):

sum the X distance and the y distance D=sqr(X^2+Y^2)

OpenStudy (anonymous):

sum X= 0 + 1.8*cos(45)-0.9*cos(30) sum Y= 3.8+1.8*sin(45)-0.9sin(30)

OpenStudy (anonymous):

cos (45)=sin(45)=sqr(2)/2 sin(30)=1/2 cos(30)=sqr(3)/2

OpenStudy (anonymous):

the answer must be in the choices: a. 4.64 b.6.46 c.78.6 d. 67.8